3. Scherz section 2.15 (p.25): another small errorplease show work
Answer #1

P= W/t

P=Vi , so

w=\int iV dt

voltage across inductor is given by V=L \frac{di}{dt}

subsitute V in W, then we get

W=\int iL \frac{di}{dt} dt

{\color{Magenta} W=\int L* i *di}

W=\frac{1}{2} Li^{2}